OpenStudy (anonymous):

find the exact value of sin(θ/2) if sinθ=-(√3/3), (3π/2)<θ<2π

OpenStudy (anonymous):

x=sin-1 (-√3/3) then multipy this value by 2

OpenStudy (anonymous):

are you saying multiple the inverse sign of (negative root 3 over 3) by 2?

OpenStudy (anonymous):

yes -√3/3 is also -1/√3

OpenStudy (anonymous):

But to find the exact value you would have to use the Half-angle formula right? \[\sin (\alpha \div2) = \sqrt{(1-\cos \alpha)\div 2}\]